# "Which LLM should I use?": Evaluating LLMs for tasks performed by Undergraduate Computer Science Students in India

### v1

@@ 3.3. Technical Interview Preparation
In this task, LLMs were prompted to generate code solutions for problems listed on LeetCode. We selected $150$ practice questions from the roadmap provided by NeetCode(Nee, 2024), a popular online resource for systematically practicing problems listed on LeetCode. These questions covered diverse topics with varying levels of difficulty such as hashing, linked list, dynamic programming, graphs, trees etc. Here is an example prompt which was provided to each LLM:
@@
The generated solutions were evaluated by submitting them to the Leetcode platform. If the submitted solution did not pass all the test cases, the information was fed back into the LLM to improve the solution. The process was repeated until it passed all the test cases or for a maximum of four attempts. At the end of this process, if the solution was failing any test, the solution was marked as incorrect. Else, the solution was marked as correct.
@@ 3.4. Learning New Concepts and Frameworks
Each LLM-generated response was rated on a scale of 1 to 10 (10 being excellent) based on factors such as depth of understanding, clarity and coherence, real-world applications, practical relevance & up-to-date academic/industrial information.
@@ 3.5. Writing Emails
This task involved testing LLMs for their capability to generate content for writing emails as emails have become a one of the most important ways of communication between students and instructors. Recognizing the pivot role of email as a communicating interface, this task evaluates LLMs on 35 prompts (our, 2024d) depicting different scenarios which an undergraduate computer science student may face at school. The prompts cover a wide spectrum of communication needs like communicating with faculty regarding coursework, feedback on projects, deadline extensions and grading, collaborating and networking with peers and personalities in academia and industry, administrative and academic queries for medical leave, scholarship, credit transfer, course and event proposals, and for personal support like forming study groups and accommodating special needs. An example prompt provided to the LLM is as follows:
"""As an undergraduate computer science student, draft emails on the topic that will be provided: Write an email to the academic department that you are unwell and cannot take the upcoming exam due to a sudden illness, providing medical documentation asking for medical leave and fixing a date to accommodate your recovery """
Each response from LLM was rated on a scale of 1 - 10 (with 10 being excellent) on multiple parameters including clarity, delivery, relevance, and tone.
@@ 4.1. Quantitative Analysis
Table 2 shows the overall quantitative results for all the tasks. For each task in this table, we are computing the average rating or average score for each LLM. We find that different LLMs are best suited for different tasks and there is no single LLM which performs well for all kinds of tasks. Microsoft Copilot performs the best for the task of code explanation and documentation, solving theoretical and humanities assignments. GitHub Copilot Chat performs the best for programming assignments and technical interview preparation as it is specifically trained for programming-related tasks. ChatGPT performs best for writing emails as well as generating solutions for technical intereview questions (tied with GitHub Copilot Chat). Lastly, Google Bard is the best performer for explaining new concepts and frameworks.
@@
Code Explaination and Documentation
8.33
0.52
7.17
2.56
9.67
0.52
8.17
1.17
Microsoft Copilot
Theoretical Assignments
7.71
1.8
2.71
2.56
8.14
1.07
6.71
1.89
Microsoft Copilot
Programming Assignments
3
1.93
5.38
2.39
2.5
2.27
7.13
1.73
Github Copilot Chat
Humanities Assignments
4.83
3.19
4.5
3.02
6.5
2.43
Microsoft Copilot
Learning new concepts and frameworks
5.33
3.26
7.75
2.93
5.67
3.08
Google Bard
Technical Interview preparation
9.65
5.73
7.69
9.65
ChatGPT and GitHub Copilot Chat
Writing Emails
8.29
1.15
6.06
1.33
4.6
2.37
3.83
2.7
ChatGPT
Table 2. Performance of Large Language Models in various tasks
@@ 4.2.1. Code Explanation and Documentation
We find that different LLMs have different proficiencies for generating code explanation and comments for different programming languages. For instance, Google Bard generates well-presented explanations with examples for SQL queries. On the other hand, it struggled with generating comments for Java and C/C++ programming languages, probably because these languaged had long code snippets. It’s responses included incomplete code snippets which did not improve even with multiple attempts. Microsoft Copilot generated very detailed code explanations and comments for all the programming languages and additionally provided relevant resources from the web. It was also able to provide mathematical and logical reasoning wherever required. ChatGPT, performed similar to Microsoft Copilot but lagged behind in providing mathematical and logical reasoning for some code snippets. Github Copilot also performed well in this task but lacked detailed explanation for code snippets where longer explanations in natural language were expected.
@@ 4.2.2. Class assignments
Programming Assignments. GitHub Copilot Chat demonstrated stronger performance among all the LLMs, showcasing its comprehensive knowledge and proficiency in the topics from the assignment set. The LLM accurately comprehended the assignment requirement, and the responses were aligned with the question demand, which significantly aided the assignment completion. It showed consistency and reliability in coding-related exercises that require a good knowledge of systems programming and C language. Bard responses’ were moderate in quality. The responses contained basic outlines and details that are good to be built upon by a student taking Bard’s help in assignments. At the same time, Bard diverted with the question deliverables on complex tasks like creating a custom syscall and versions of dining philosophers, which further damped the helpfulness of its responses. On the other hand, ChatGPT displayed basic understanding of the tasks and interpreted the questions well, providing a high level overview of the solution but its responses lacked details of implementation and specific linux commands. Microsoft Copilot was least performant in this task. It’s limitations were apparent in providing answers to exercises that required knowledge of intricate subjects such as thread scheduling, inter-process communication, syscalls, kernel modules, and versions of the dining philosophers problem. It’s responses were vague and lacked requisite depth, offering limited assistance towards finding the solution.
Theoretical Assignments. Microsoft Copilot demonstrated strong performance across the set of problems we used for evaluation. It was consistent in providing strong responses with correct algorithms, pseudo code and time complexities. GitHub Copilot Chat was competent to provide the correct pseudocode, but it was not giving time complexities at majority of the times. Despite the lack of time complexity information, its responses were found to be fairly helpful for completing the assignments. Similar challenges were found with ChatGPT as it was also giving partially correct answers.
Google Bard faced challenges in delivering correct results failing to understand the intricacies of the problem. There were instances when the output algorithm contained different procedures than those required by the question. This compromised the completeness and correctness of the solutions and gave wrong time complexities to solve the problem. For instance, a question where Bard diverted from the correct approach to solve the problem was as follows:
@@
The response given by Bard to the above problem describes a backtracking approach with given time complexity of $O\left(n^{4}\right)$, whereas the correct solution requires a Divide and Conquer paradigm based approach with complexity of $O\left(n^{2}\right)$. In conclusion, the LLMs exhibited diverse patterns while tending to non-trivial questions. Microsoft Copilot delivered robust and consistently good results, while ChatGPT and GitHub Copilot Chat were vulnerable to giving only partially correct algorithms. Notably, Bard’s results were the least helpful among all the models that were evaluated.
Humanities Assignments. Across the multiple dimensions tested, ChatGPT demonstrated a propensity for delivering responses with a bookish quality, lacking the critical thinking required in several questions. In contrast, Microsoft Copilot emerged as the standout performer, particularly in debate and discussion on sustainable development goals, showcasing a commendable ability to navigate and debate from multiple perspectives, as well as providing the most current and relevant updates on recent affairs. The model excelled in sociology theorems and rising innovations by delivering informative responses with critical solutions and intuitive thinking. On the other hand, Google Bard’s performance varied across tasks, exhibiting strengths in real-time solutions but falling short in areas such as debate. In summary, while ChatGPT struggled to infuse critical thinking, Microsoft Copilot emerged as the most adept in handling the complexity of such topics, offering up-to-date insights and creative solutions, thereby positioning itself as a valuable resource for students and researchers in Humanities.
@@ 4.2.3. Technical Interview Preparation
While testing LLMs on beginner-level topics like arrays, searching, stacks, pointers, greedy and linked list, ChatGPT, Microsoft Copilot, and Copilot Chat demonstrated commendable accuracy across difficult levels of questions, showcasing their competence in foundational concepts. However, LLMs showed varying trends when subjected to questions involving intermediate to advanced concepts like trees, tries, graphs, dynamic programming, backtracking and heaps. Microsoft Copilot showed inconsistent performance with its responses fluctuating between being unable to answer and giving correct solutions in one go. Google Bard consistently fell short in providing accurate results in the first attempt across all the levels of topic complexity and question difficulties and after the correction of response the accuracy was still less as compared to its counterparts. GitHub Copilot Chat and ChatGPT were the best performers among all the LLMs, with high accuracy through the tests. GitHub Copilot Chat had remarkable consistency in providing correct solutions across difficulty levels with soundness in solving algorithm problems, while ChatGPT showcased the most interpretable responses. ChatGPT was additionally giving easy-to-understand explanations in addition to the correct code while GitHub Copilot Chat only generated correct code with minimal explanation and comments. An example where Github Copilot Chat and ChatGPT were correct with their responses while Microsoft Copilot and Bard struggled to give correct solutions even after the fourth attempt is N-Queens Problem, which is tagged as Hard on LeetCode website from the topic backtracking.
@@ 4.2.4. Learning new Concepts and Frameworks
Concept tasks, designed to assess the models’ proficiency in grasping and articulating new ideas, showcased ChatGPT’s notable strengths in specific areas. Despite its inclination towards a bookish quality and a propensity to lack critical thinking in intricate questions, ChatGPT demonstrated efficacy in tasks such as summarizing papers and finding similarities and differences among them. Moving to framework tasks, which gauge the models’ aptitude in suggesting, explaining, and providing resources for understanding frameworks, Microsoft Copilot demonstrated its overall robust performance. Some tasks required the models to suggest frameworks and compare them, with Microsoft Copilot excelling in clarity and coherence, real-world applications, and practical relevance & up-to-date academic/industrial information. The assessment further delved into the models’ responses to various tasks such as learning a new language by chatting with it, responding to easy, medium, and hard questions, and correcting errors of a programmer. Microsoft Copilot consistently demonstrated a commendable blend of clarity, depth of understanding, and real-world applicability in these diverse tasks, reinforcing its versatile capabilities. In conclusion, while ChatGPT and Bard exhibit diverse proficiency, Microsoft Copilot emerged as the best performer, showcasing consistent strengths in critical thinking, real-world relevance, and depth of understanding.
@@ 4.2.5. Writing Emails
Microsoft Copilot and Google Bard displayed a mix of responses varying between weak responses and strong responses, with some vague responses lacking the required details as mentioned in the prompts. Microsoft Copilot’s responses occasionally had incomplete sentences, affecting the coherence of the responses. GitHub Copilot Chat was unable to respond to demanding prompts that involved creative composition and generating details and ideas as per the prompt. Responses were inconsistent and lacked relevance in comparison to other LLMs. This is expected as GitHub is specifically aimed at assisting in programming-related tasks. For example, when prompted with the following prompt: """Write an email to your university’s accessibility services office to request accommodations for a disability that affects your ability to complete coursework.""", GitHub Copilot Chat denied responding to this prompt while Microsoft Copilot only gave a high level outline of the email. On the other hand, ChatGPT consistently delivered moderate to high quality responses across all the prompts. It displayed good clarity, delivery, relevance and tone, effectively addressing the demand of the prompts to help undergraduate computer science students in various scenarios. ChatGPT’s responses were easily distinguishable from other LLMs when the prompt required the LLM’s response to be sophisticated and showcase its knowledge of subject matter and effectively articulate it into an email.

### v2

@@ 3.3. Technical Interview Preparation
In this task, LLMs were prompted to generate code solutions for problems listed on LeetCode.com. We selected $150$ practice questions from the roadmap provided by NeetCode(Nee, 2024), a popular online resource for practicing problems listed on LeetCode to help students prepare for technical interviews. Leetcode tags the problems in three difficulty levels namely easy, medium, and difficult. These questions covered diverse topics such as dynamic programming, greedy approach, divide and conquer approach, hashing, sliding window approach, stacks, linked lists, graphs, trees, etc. Here is an example prompt which was provided to each LLM:
@@
The generated solutions were evaluated by submitting them to the Leetcode platform. If the submitted solution did not pass all the test cases on Leetcode.com, the information of the failing testcase or the error that was generated was fed back into the LLM to improve the solution. This iterative process continued until the solution successfully passed all the test cases or until a maximum of 3 correction attempts were made. At the end of this process, if the solution was failing any testcase, it was marked as incorrect. Else, the solution was marked as correct.
@@ 3.4. Learning New Concepts and Frameworks
The responses were rated on a scale of 1 - 10. All the parameters were given equal weightage while assigning the rating to responses. Metrics for evaluation of the task are defined as follows:
  • Depth of understanding: How well the responses are able to comprehend computer science concepts, including intricacies and details of different frameworks.
  • Clarity and Coherence: How easily interpretable were the responses and how logically did the LLM convey information, ensuring that the explanations and guidance provided were easy to follow and free from ambiguity.
  • Real-world Applications and Practical Relevance: How appropriately do the LLM relate the concepts and how applicable is the information provided by LLM to real-world scenarios and practical applications in computer science.
  • Up-to-date Academic/Industrial Information: The inclusion of current and relevant information from academic sources or industry standards.
@@ 3.5. Writing Emails
This task evaluated the LLMs’ capability to generate content for writing emails, as email is one of the most crucial communication mediums for students in both academic and professional settings. This task evaluates LLMs on 35 prompts (our, 2024d) depicting different scenarios that an undergraduate computer science student may face at both university and industry. The prompts cover a wide spectrum of communication needs like communicating with faculty regarding coursework, feedback on projects, deadline extensions, doubts in grading, collaborating and networking with peers and personalities in academia and industry, administrative and academic queries for medical leave, scholarship, credit transfer, course and event proposals, and for personal support like forming study groups and accommodating special needs. An example prompt provided to the LLM is as follows:
"""As an undergraduate computer science student, draft emails on the topic that will be provided: Write an email to the academic department that you are unwell and cannot take the upcoming exam due to a sudden illness, providing medical documentation asking for medical leave and fixing a date to accommodate your recovery """ The responses were rated on a scale of 1 - 10. All the parameters were given equal weightage while assigning the rating to responses. Metrics for evaluation of the task are defined as follows:
  • Clarity: How well the responses conveyed information in an unambiguous and interpretable manner. An example of poor and good clarity in the context of writing emails by an undergraduate student:
    Poor Clarity: "I need help with the topic we learned in last lecture."
    Good Clarity: "I need help understanding the concept of ’Network Protocols’ that we covered in the last Computer Networks lecture."
  • Delivery: How effectively was the information presented. Examples of text with poor delivery and good delivery are given below.
    Poor Delivery: "Attached with the email is assignment. Let me know of any issues."
    Good Delivery: "I’ve attached my Assignment-2 for Data Structures and Algorithms course. If there are any issues or if you need additional information, please let me know."
  • Relevance: How appropriately does the response address the details in the prompt, avoiding unnecessary information. An example is as follows:
    Poor Relevance: "I enjoyed watching the F1 Grand Prix yesterday. Also, I can’t make it to tomorrow’s class."
    Good Relevance: "Due to a personal commitment, I won’t be able to attend tomorrow’s ‘Machine Learning’ class. Could you please share any notes or updates that I may miss?"
  • Tone: How is the style of expression in the language of response, affecting the interpersonal impact of communication. Example of a poor tone in email writing and a good tone in email writing:
    Poor Tone: "You must send me the work by noon."
    Good Tone: "Could you please send me the project report by noon?"
@@ 4.1. Quantitative Analysis
Table 3 shows the overall quantitative results for all the tasks. The table contains scores calculated using the average across all the ratings received for all the responses per task, along with standard deviation showcasing the fluctuation in response quality based on respective metrics of tasks. We find that different LLMs are best suited for different tasks and no single LLM performs well for all kinds of tasks. Microsoft Copilot performs the best for the task of code explanation and documentation, solving theoretical and humanities assignments. GitHub Copilot Chat performs the best for programming assignments and technical interview preparation as it is specifically trained for programming-related tasks. ChatGPT performs best for writing emails as well as generating solutions for technical intereview questions (tied with GitHub Copilot Chat). Lastly, Google Bard is the best performer for explaining new concepts and frameworks.
Table 2 presents the accuracy percentages of different LLMs for Technical Interview Preparation, calculated based on the number of correct solutions. In this task, Github Copilot Chat and ChatGPT achieve the highest accuracy of 96.5% while Google Bard has the least accuracy of 57.3%.
LLM Easy Questions Medium Questions Hard Questions Total Accuracy %
First Attempt Correct Subsequent Attempt Correct Total questions in set Accuracy % First Attempt Correct Subsequent Attempt Correct Total questions in set Accuracy % First Attempt Correct Subsequent Attempt Correct Total questions in set Accuracy %
ChatGPT 27 1 28 100 87 3 93 96.77 19 1 22 90.91 96.5
Google Bard 13 10 28 82.14 30 26 93 60.22 0 3 22 13.64 57.34
Microsoft Copilot 27 0 28 96.43 68 1 93 74.19 13 1 22 63.64 76.92
GitHub Copilot Chat 27 1 28 100 87 2 93 95.7 19 2 22 95.45 96.5
Table 2. Accuracy Percentage of Large Language Models in Solving Questions on Leetcode.com
@@
Code Explanation and Documentation 8.33 0.52 7.17 2.56 9.67 0.52 8.17 1.17 Microsoft Copilot
Theoretical Assignments 7.71 1.8 2.71 2.56 8.14 1.07 6.71 1.89 Microsoft Copilot
Programming Assignments 3 1.93 5.38 2.39 2.5 2.27 7.13 1.73 Github Copilot Chat
Humanities Assignments 4.83 3.19 4.5 3.02 6.5 2.43 0.75 1.8 Microsoft Copilot
Learning new concepts and frameworks 5.33 3.26 7.75 2.93 5.67 3.08 5 3.67 Google Bard
Writing Emails 8.29 1.15 6.06 1.33 4.6 2.37 3.83 2.7 ChatGPT
Table 3. Performance of Large Language Models in various tasks | Scale: 1 - 10 (Higher is better)
@@ 4.2.1. Code Explanation and Documentation
We find that different LLMs have different proficiency for generating code explanations and comments for different programming languages. For instance, Google Bard generates well-presented explanations with examples for SQL queries. On the other hand, it struggled with generating comments for Java and C/C++ programming languages, which was likely because of the length of code snippets involved in the evaluation. Its responses included incomplete code snippets that failed to meet the required standards, despite receiving up to five rounds of feedback from the evaluator. It performed relatively better in Python Code. Microsoft Copilot generated very detailed code explanations and comments for all the programming languages and additionally provided relevant resources from the web to search for additional information about the topic. It was also able to provide logical reasoning wherever required. For instance, it was correctly able to figure out why we can set A + $A^{T}$ to 0 in the case of a skew-symmetric matrix, which Bard and ChatGPT could not. ChatGPT performed similarly to Microsoft Copilot but lagged behind in providing logical reasoning in 2 out of 10 prompts. GitHub Copilot Chat demonstrated good logical reasoning abilities in this task and was able to figure out the A + $A^{T}$ case as described above. However, it encountered difficulties when asked to provide in-depth explanations for code snippets, particularly when those explanations required longer and more detailed responses in plain English.
@@ 4.2.2. Class assignments
Programming Assignments. GitHub Copilot Chat demonstrated stronger performance among all the LLMs, showcasing its comprehensive knowledge and proficiency in the topics from the assignment set. The LLM accurately comprehended the assignment requirement, and the responses aligned with the question demand, significantly aiding the assignment completion. It showed consistency and reliability in coding-related exercises requiring good systems programming and C language knowledge. Bard’s responses’ were moderate in quality. The responses contained basic outlines and details that are good to be built upon by a student taking Bard’s help in assignments. At the same time, Bard diverted with the question deliverables on complex tasks like creating a custom syscall and versions of dining philosophers problem, which further damped the helpfulness of its responses. On the other hand, ChatGPT displayed a basic understanding of the tasks and interpreted the questions well, providing a high-level overview of the solution but its responses lacked details of implementation and specific Linux commands. Microsoft Copilot was the worst performer in this task. Its limitations were apparent in providing answers to exercises that required knowledge of intricate subjects such as thread scheduling, inter-process communication, syscalls, kernel modules, and versions of the dining philosophers problem. Its responses were vague and lacked requisite depth, offering limited assistance in finding the solution.
Theoretical Assignments. Microsoft Copilot demonstrated strong performance across the set of problems we used for evaluation. It was consistent in providing strong responses with correct algorithms, pseudo-code, and time complexities. GitHub Copilot Chat was competent to provide the correct pseudo-code, but it did not give the time complexity of the solution in six out of seven experiments. Despite the lack of time complexity information, its responses were found to be fairly helpful for completing the assignments. Similar challenges were found with ChatGPT as it was also giving partially correct answers. Google Bard faced challenges in delivering correct results failing to understand the intricacies of the problem. There were instances when the output algorithm contained different procedures than those required by the question. This compromised the completeness and correctness of the solutions and gave wrong time complexities to solve the problem. For instance, a question where Bard diverted from the correct approach to solve the problem was as follows:
@@
The response given by Bard to the above problem describes a backtracking approach with a given time complexity of $O\left(n^{4}\right)$, whereas the correct solution requires a Divide and Conquer paradigm based approach with the complexity of $O\left(n^{2}\right)$. In conclusion, the LLMs exhibited diverse patterns while tending to non-trivial questions. Microsoft Copilot delivered robust and consistently good results, while ChatGPT and GitHub Copilot Chat were vulnerable to giving only partially correct algorithms. Notably, Bard’s results were the least helpful among all the models that were evaluated.
Humanities Assignments. Across the multiple dimensions tested, ChatGPT often delivered responses with a bookish quality, lacking the critical thinking required in several questions. In contrast, Microsoft Copilot emerged as the standout performer, particularly in debate and discussion on sustainable development goals, showcasing a commendable ability to navigate and debate from multiple perspectives, as well as providing the most relevant and recent updates on current affairs. The model excelled in sociology theorems and rising innovations by delivering informative responses with critical solutions and intuitive thinking. On the other hand, Google Bard’s performance varied across tasks, exhibiting strengths in real-time solutions but falling short in areas such as debate. GitHub Copilot’s assistance is limited to programming-related inquiries only, as it lacked knowledge of humanities assignments, stating, "I can only assist with programming-related questions" and "My expertise is strictly limited to software development topics." As a result, it was not effective in providing assistance for humanities assignments. In summary, while ChatGPT struggled to infuse critical thinking, Microsoft Copilot emerged as the most adept in handling the complexity of such topics, offering up-to-date insights and creative solutions, thereby positioning itself as a valuable resource for students and researchers in Humanities.
@@ 4.2.3. Technical Interview Preparation
While testing LLMs on beginner-level topics like arrays, searching, stacks, pointers, greedy and linked lists, ChatGPT, Microsoft Copilot, and Copilot Chat demonstrated commendable accuracy across difficult levels of questions, showcasing their competence in foundational concepts. However, LLMs showed varying trends when subjected to questions involving intermediate to advanced concepts like trees, tries, graphs, dynamic programming, backtracking, and heaps. Microsoft Copilot showed inconsistent performance with its responses fluctuating between being unable to answer and giving correct solutions in one go. Google Bard consistently fell short in providing accurate results in the first attempt across all the levels of topic complexity and question difficulties and even after the correction attempts of response the accuracy was still less as compared to its counterparts. GitHub Copilot Chat and ChatGPT were the best performers among all the LLMs, with high accuracy through the tests. GitHub Copilot Chat consistently provided correct solutions across difficulty levels with soundness in solving algorithm problems, while ChatGPT showcased the most interpretable responses. ChatGPT was additionally giving easy-to-understand explanations to the solution and its approach in addition to the correct code while GitHub Copilot Chat only generated correct code with minimal explanation and comments. An example where Github Copilot Chat and ChatGPT were correct with their responses while Microsoft Copilot and Bard struggled to give correct solutions even after the third correction attempt is N-Queens Problem, which is tagged as Hard on LeetCode website from the topic backtracking.
@@ 4.2.4. Learning new Concepts and Frameworks
Learning new concept tasks assessed how well the LLMs could help users learn new concepts by providing clear explanations and resolving any follow-up questions or doubts. ChatGPT exhibited notable strengths in certain areas, such as summarizing papers and identifying similarities and differences among them. However, it sometimes lacked critical thinking for complex questions and displayed a tendency towards a bookish style.
We also evaluated the LLM’s ability to suggest programming frameworks, explain them, and provide resources for understanding them. Additionally, we assessed their performance in assisting with learning a new programming language. Google Bard demonstrated strong overall performance. For tasks that needed suggestions of references or clear and detailed information, Google Bard shined by also providing explanatory reasons. It also offers up-to-date academic/industrial information because it can access the web, which is particularly beneficial for programming frameworks that are frequently updated. The responses provided by Microsoft Copilot were clear and coherent and had up-to-date information just like Google Bard, but the answers lacked depth of understanding and were a bit too concise. In terms of up-to-date academic/industrial information, Google Bard has an advantage over Microsoft Copilot as it uses Google’s superior search engine (Lewandowski, 2014) for Information Retrieval, while Copilot relies on Bing’s less comprehensive search results.
In summary, ChatGPT is better at explaining new concepts in an easy-to-understand way for learning purposes. GitHub Copilot Chat struggles to give detailed explanations in plain English, specifically for non-coding tasks. However, it excels at providing responses that need technical details, such as subtle details about programming frameworks like Django. Microsoft Copilot provided current and relevant information with useful resources but lacked clarity and detail. Among the LLMs evaluated, Google’s Bard demonstrated the best performance on this task. Bard consistently provided clear responses exhibiting strong logical reasoning abilities and incorporating up-to-date academic/industrial knowledge.
@@ 4.2.5. Writing Emails
Microsoft Copilot and Google Bard displayed a mix of weak and strong responses, with vagueness in responses lacking the required details as mentioned in the prompts. In nine out of thirty-five scenarios, Microsoft Copilot’s responses had incomplete sentences, affecting their coherence. GitHub Copilot Chat was unable to respond to demanding prompts that involved creative composition and generating details and ideas as per the prompt. Responses were inconsistent and lacked relevance in comparison to other LLMs. This is expected as GitHub is specifically aimed at assisting in programming-related tasks. For example, when prompted with the following prompt: """Write an email to your university’s accessibility services office to request accommodations for a disability that affects your ability to complete coursework.""", GitHub Copilot Chat denied responding to this prompt while Microsoft Copilot only gave a high-level outline of the email. On the other hand, ChatGPT consistently delivered moderate to high-quality responses across all the prompts. It displayed good clarity, delivery, relevance and tone, effectively addressing the demand of the prompts to help undergraduate computer science students in various scenarios. ChatGPT’s responses were easily distinguishable from other LLMs when the prompt required the LLM’s response to be sophisticated, showcase its knowledge of the subject matter and effectively articulate it into an email.